During 2017, Vaughn Manufacturing expected Job No. 26 to cost $300000 of overhead, $500000 of materials, and $200000 in labor. Vaughn applied overhead based on direct labor cost. Actual production required an overhead cost of $360000, $640000 in materials used, and $300000 in labor. All of the goods were completed. What amount was transferred to Finished Goods?

$1300000
$1240000
$1390000
$1000000
Answer #1

Overhead rate = Estimated overhead/Estimated labor cost

= 300,000/200,000 = 150% of dl cost

Amount transferred to Finished goods

= Direct Materials + Direct Labor + Overhead

= 640,000 + 300,000 + (300,000*150%)

= 1,390,000

Option C